OpenStudy (anonymous):

Determine the zeros of the function x^2 − 25 = 0.

OpenStudy (campbell_st):

this is the difference of 2 squares \[(x -a)(x+a) = x^2 - a^2\] match your equation to the standard form above.. and you should get the answer

OpenStudy (anonymous):

Factor x^2-25

OpenStudy (anonymous):

I got (x-5)(x-5)=0 but that's not it

OpenStudy (campbell_st):

nearly (x -5)(x + 5) = 0 now you need to solve the binomial factors x - 5 = 0 x + 5 = 0 there will be 2 values of x...

OpenStudy (anonymous):

oh is it 0?

OpenStudy (campbell_st):

nope... find the values of x that make x - 5 = 0 as well as x + 5 = 0

OpenStudy (anonymous):

-5 and +5

OpenStudy (anonymous):

?

OpenStudy (campbell_st):

thats it... well done

OpenStudy (anonymous):

Can you help me more

OpenStudy (anonymous):

Determine the zeros of the function 7x2 + 9x = 0.

OpenStudy (anonymous):

*Determine the zeros of the function 7x^2 + 9x = 0.

OpenStudy (campbell_st):

ok... so you need to factor the equation... what is a common factor between 7x^2 and 9x ?

OpenStudy (anonymous):

1?

OpenStudy (campbell_st):

well what about a letter factor...?

OpenStudy (anonymous):

oh x hahaha

OpenStudy (campbell_st):

great so the factored form is x(7x + 9) = 0 now if either factor is zero, then the whole equation would equal zero... so 1 solution is x = 0 the other solution is 7x + 9 = 0 now just solve the equation for x

OpenStudy (anonymous):

9/7 ? is that the other answer

OpenStudy (anonymous):

:(

OpenStudy (campbell_st):

well its -9/7... so the 2 solutions are x = 0 and x = -9/7 substitute either value into the equation and you'd have zero as the answer...

OpenStudy (anonymous):

okay thank you can you keep helping me?

OpenStudy (anonymous):

This

OpenStudy (campbell_st):

ok... this is the graphical solution... the zeros... are the the points where y = 0.... find the corresponding x values... which means, read off the x values where the curve cuts the x-axis.... hope it all helps..

OpenStudy (anonymous):

I don't know how to do that

OpenStudy (campbell_st):

the x- axis is the horizontal ... where does the curve cut the horizontal axis... there are 2 points